The sum of three consecutive intergers is 7 less than twice the first. What are the intergers?
Answer by JulietG(1812)

A + B + C = 2A - 7
B = A + 1
C = A + 2
Substitute
A + (A+1) + (A+2) = 2A-7
Add
3A + 3 = 2A -7
subtract 2A from each side
A + 3 = -7
Subtract 3 from each side
A = -10
Numbers are -10, -9, -8
.
Add them together to get -27
-27 is 7 less than twice -10.